two typists working together can type a report in 2.4 hrs. if either had to type the report all alone, it would take the 1st typist 2 hrs longer than the 2nd typist. how long would it take for each typist to type the said report?
Answer by htmentor(1343)

Let t = the number of hours the 2nd typist takes working alone
So the rate of the 2nd typist is 1/t
Since the 1st typist takes t+2 hours, their rate is 1/(t+2)
Working together, their combined rate is 1/t + 1/(t+2) = 1/2.4
Solve for t:
(2t+2)/(t^2+2t) = 1/2.4
t^2 -2.8t - 4.8 = 0
Factor:
(t+1.2)(t-4) = 0
We take the positive solution, t = 4
Thus the 1st typist takes 6 hours, and the 2nd typist takes 4 hours